Axolotl:

Please help with Math question! Thanks :)))) Which of the following tables shows the correct steps to transform x2 + 10x + 24 = 0 into the form (x − p)2 = q? [p and q are integers] Step 1 x2 + 10x + 24 − 1 = 0 − 1 Step 2 x2 + 10x + 23 = −1 Step 3 (x + 5)2 = −1 Step 1 x2 + 10x + 24 − 2 = 0 − 2 Step 2 x2 + 10x + 22 = −2 Step 3 (x + 5)2 = −2 Step 1 x2 + 10x + 24 + 2 = 0 + 2 Step 2 x2 + 10x + 26 = 2 Step 3 (x + 5)2 = 2 Step 1 x2 + 10x + 24 + 1 = 0 + 1 Step 2 x2 + 10x + 25 = 1 Step 3 (x + 5)2 = 1

Vocaloid:

The process of turning a quadratic equation into a perfect square is called completing the square x2 + 10x + 24 = 0 Take the x term (10x). The coefficient is 10. Take this number, divide it by 2, square it. 10/2 = 5, 5^2 = 25 So x2 + 10x + 25 would be a perfect square trinomial. x2 + 10x + 25 = (x + 5)^2 However, we have x^2 + 10x + 24 instead of + 25. So we add 1 to both sides to get x^2 + 10x + 24 + 1 = 0 + 1 And rewrite the expression from there, converting the left side to (x+5)^2 and simplifying the right side to 1